# Home Swapping – An Innovative Approach to Reduce Traffic Congestion and Carbon Emissions

### v1

@@ VI Greedy Home Swapping
Change in Departure Times. The empirical distribution of commuters’ departure time is shown in Figure S2A. As we can see, commuters departure times peak between 7-8 AM, with the highest number of departures at 7:30 AM. It is stipulated that if the morning arrival time to the workplace and the evening departure time from the workplace remain unchanged, the new morning departure time from their new home and the evening new arrival time back to their new home after GHS would shift and can be deduced from navigation data.
The distribution of the commuters’ new departure time from their new home after GHS is shown in Figure S3A. The results revealed that the departure time becomes more dispersed compared to that in the empirical data, with fewer residents departing during peak time (7:30 AM). The distribution of the new departure times after SDGHS with medium tolerance is shown in Figure S3B, which is similar to that after GHS, but the number of commuters who use motor vehicles increases, resulting in a lower reduction in commuting distance, i.e. $\Delta d$.
Supplementary Figure S3. The Distribution of Home-Departure Times During the Morning Rush Hour. Panel (A) shows the departure time after GHS; Panel (B) shows the departure time after SDGHS under medium tolerance.
@@
Supplementary Figure S4. The congestion coefficient obtained based on the empirical data and after RHS and GHS. The top three maps show the congestion coefficient during the morning peak hours between 7-8 AM; the middle three maps show the congestion coefficient during the morning peak hours between 8-9 AM; the bottom three maps show the congestion coefficient during the evening peak hours between 5-6 PM.
@@
Supplementary Figure S6. The number of congested, slow-moving, and free-flowing road segments during the morning rush hour between 7-8 AM.
Since the average congestion coefficients $I_{t}$ after GHS under different factors do not vary significantly, we further analyze the changes in the number of congested roads with $c_{\alpha}\geq 1$, slow-moving roads with $10.4\leq c_{\alpha}<$ 1, and free-flowing roads with $c_{\alpha}<0.4$ during the morning rush hour between 7-8 AM under different restrictive factors of GHS, as shown in Figure S6. After GHS, there is a noticeable decrease in congested and slow-moving road segments, while free-flowing road segments significantly increase. By considering only the amenity accessibility (a-GHS), housing price (p-GHS) and the distance to the city center (c-GHS), as well as all the three socio-demographic factors (SDGHS), the average congestion coefficient $I_{t}$ gradually increases, the number of severely congested and slow-moving road segments gradually increases, while the number of free-flowing road segments decreases. Although the number of congested roads in a-GHS is lower than that after GHS, slow-moving segments increase, and free-flowing road segments decrease. Additionally, $I_{t}$ after p-GHS is similar to that after a-GHS, but with a higher number of congested and slow-moving segments.
Considering dual socio-demographic factors (medium tolerance) during home-swapping. We also consider $\Delta d$ by considering two out of the three socio-demographic factors in GHS, as shown in Figure S7. In this discussion, we use the medium tolerance SDGHS discussed in Figure 4 of the main text. The reduction in commuting distance, i.e. $\Delta d$ under (1) the joint consideration of the distance to the city center and amenity accessibility (c&a-GHS) is 23.6%, (2) the joint consideration of of the distance to the city center and housing price (c&p-GHS) is 26.6%, and (3) the joint consideration of housing price and amenity accessibility (p&a-GHS) reaches 36.29%. The results show that the distance from the city center has a significant impact on the benefit brought by GHS, followed by amenity accessibility, and finally the housing price.
@@
The Impact of the Willingness to Swap Home. In GHS, we assume that all households are willing to swap their home as long as the total commuting distance of the two swapping households decreases, even at the expense of an increase in the commuting distance of one of them. Here, we assume that the households with an increase of commuting distance has an $x\%$ willingness to swap but $1-x\%$ to reject.
@@
Supplementary Figure S8. $\Delta d$ after GHS or after SDGHS (medium tolerance) considering households’ willingness to change residences.
@@
As different households have different levels of willingness to swap homes, $\Delta d$ after GHS method with different willingness $x\%$ is shown by the blue line in Figure S8. The results show that $\Delta d=40.9\%$ when households do not swap residences if the average commuting distance of one household increases. As the willingness to swap increases from 0% to 25%, there is a noticeable increase in $\Delta d$. However, from a willingness of 25% onwards, the increase in $\Delta d$ slows down as the willingness of swapping $x\%$ increases.
@@
Moreover, we consider households willingness to swap home in SDGHS under medium tolerance and analyzes $\Delta d$ after home swapping under different willingness $x\%$, as shown by the purple line in Figure S8. The results find that as the $x\%$increases, $\Delta d$ also increases. When $x=0$, $\Delta d$ is 9.6%; when $x>=25$, there is no significant increase in $\Delta d$.
VII The Reduction of Carbon Emissions.
@@
Supplementary Figure S9. The Road carbon emissions obtained based on the empirical data and after RHS and GHS. The top three diagrams show the CO2 emissions on the roads during the morning peak hours between 7-8 AM; the middle three diagrams show the CO2 emissions during the morning peak hours between 8-9 AM; the bottom three diagrams show the carbon emissions during the evening peak hours between 5-6 PM.
@@
We computed and visualized the carbon dioxide emissions of each road segment during the morning peak hours between 7-8 AM, between 8-9 AM, and the evening peak hour between 5-6 PM based on the empirical data and after RHS and GHS , as shown in Figure S9. We found that the carbon emissions under RHS at are greater than that based on the empirical data at all time, while those after GHS are significantly less than that of the empirical data. The carbon dioxide emissions between 7-8 AM during the morning peak hours are higher than those between 5-6 PM during the evening peak hours. Additionally, after RHS, most road segments emit a large amount of CO2, whereas after GHS, only a few segments have higher emissions, and most road segments emit a small amount of CO2 (road segments emitting 0.4 tons of carbon are marked in red, and those emitting between 0.2 and 0.4 tons are in yellow, and those emitting less than 0.2 tons are in green).
@@
The reduction in carbon emissions after GHS, and SDGHS is shown in Figure S10 for each hour during the peak commuting hours. In the morning peak hours, the average commuting distance after home swapping become shorter for most households. After home swapping, each individual is supposed to arrive at their workplace at the same time before home swapping, which leads to a dispersed departure time (see Figure S3). It was found that the reduction in carbon emissions decreases over time. Similarly for the evening peak hours, where each individual is supposed to leave workplace at the same time after home swapping, the shorter commuting distances lead to a gradual increase in carbon emissions reduction over time. Among all the cases, the GHS emissions reduction is the highest, and the emissions reduction of SDGHS is the lowest.
@@
VIII Adding new activity centers based on the original city center.
@@
IX Impact of Commuting Mode Shift Threshold.
@@
While we have assumed 2.5 km to be the shortest distance for commuters to travel by motor vehicles, we also examined other thresholds of assumed minimum distance at $2$, $1.5$, $1$, $0.5$, and 00 km, as shown in Figure S14. As the threshold decreases, the proportion of non-motorized vehicles shows a significant decline. Under medium tolerance with SDGHS, 9% of commuters live and work within a very short commuting distance.
@@
In addition, we also explore the impact of the assumed thresholds of the minimum distance of the use of non-motorized vehicles on the average congestion coefficient and carbon emissions. Figure S15A shows the impact of different values of thresholds on traffic congestion during the 7-8 AM peak hours. As the threshold value decrease, the number of used motor vehicles increases, the occupancy of road sections during peak hours increases, thus increasing congestion. When all commuters travel by motor vehicles, congestion also increases. By considering the three socio-demographic factors, there are fewer non-motorized vehicles, and the change in the minimum travel distance for using motor vehicles does not significantly impact traffic congestion.

### v2

@@ VI Greedy Home Swapping
Change in departure time. The empirical distribution of commuters’ departure time is shown in Figure S2A. As we can see, the departure time of commuters peaks between 7-8 AM, with the highest number of departures at 7:30 AM. It is stipulated that if the morning arrival time to the workplace and the evening departure time from the workplace remain unchanged, the new morning departure time from their new home and the evening new arrival time back to their new home after GHS would shift and can be deduced from navigation data.
The distribution of the commuters’ new departure time from their new home after GHS is shown in Figure S3A. The results revealed that the departure time becomes more dispersed compared to that in the empirical data, with fewer residents departing during peak time (7:30 AM). The distribution of the new departure time after SDGHS with medium tolerance is shown in Figure S3B, which is similar to that after GHS, but the number of commuters who use motor vehicles increases, resulting in a lower reduction in commuting distance, i.e. $\Delta d$.
Supplementary Figure S3. The Distribution of Home-Departure Time During the Morning Rush Hour. Panel (A) shows the departure time after GHS; Panel (B) shows the departure time after SDGHS under medium tolerance.
@@
Supplementary Figure S4. The congestion coefficient obtained based on the empirical data and after RHS and GHS. The top three maps show the congestion coefficient during the morning peak hour between 7-8 AM; the middle three maps show the congestion coefficient during the morning peak hour between 8-9 AM; the bottom three maps show the congestion coefficient during the evening peak hour between 5-6 PM.
@@
Supplementary Figure S6. The number of congested, slow-moving, and free-flowing road segments during the morning rush hour between 7-8 A.M.
Since the average congestion coefficients $I_{t}$ after GHS under different factors do not vary significantly, we further analyze the changes in the number of congested roads with $c_{\alpha}\geq 1$, slow-moving roads with $0.4\leq c_{\alpha}<1$, and free-flowing roads with $c_{\alpha}<0.4$ during the morning rush hour between 7-8 AM under different restrictive factors of GHS, as shown in Figure S6. After GHS, there is a noticeable decrease in congested and slow-moving road segments, while free-flowing road segments significantly increase. By considering only the amenity accessibility (a-GHS), housing price (p-GHS) and the distance to the city center (c-GHS), as well as all the three socio-demographic factors (SDGHS), the average congestion coefficient $I_{t}$ gradually increases, the number of severely congested and slow-moving road segments gradually increases, while the number of free-flowing road segments decreases. Although the number of congested roads in a-GHS is lower than that after GHS, slow-moving segments increase, and free-flowing road segments decrease. Additionally, $I_{t}$ after p-GHS is similar to that after a-GHS, but with a higher number of congested and slow-moving segments.
Considering dual socio-demographic factors (medium tolerance) during home-swapping. We also consider $\Delta d$ by considering two out of the three socio-demographic factors in GHS, as shown in Figure S7. In this discussion, we use the medium tolerance SDGHS discussed in Figure 4 of the main text. The reduction in average commuting distance, i.e. $\Delta d$ under (1) the joint consideration of the distance to the city center and amenity accessibility (c&a-GHS) is 23.6%, (2) the joint consideration of the distance to the city center and housing price (c&p-GHS) is 26.6%, and (3) the joint consideration of housing price and amenity accessibility (p&a-GHS) reaches 36.29%. The results show that the distance to the city center has a significant impact on the benefit brought by GHS, followed by amenity accessibility, and finally the housing price.
@@
The impacts of the willingness and selfishness to swap home. In GHS, we assume that all households are willing to swap their homes as long as the total commuting distance of the two swapping households decreases, even at the expense of an increase in the commuting distance of one of them. Here, we assume that the households with an increase of commuting distance has an $w\%$ willingness to swap but $1-w\%$ to reject.
@@
Supplementary Figure S8. (A) $\Delta d$ after GHS or after SDGHS (medium tolerance) considering households’ willingness to change residences. (B) $\Delta d$ after GHS or after SDGHS (medium tolerance) considering household members’ selfishness to change residences.
@@
As different households have different levels of willingness to swap homes, $\Delta d$ after GHS method with different willingness $w\%$ is shown by the blue line in Figure S8A. The results show that $\Delta d=40.9\%$ when households do not swap residences if the average commuting distance of one household increases. As the willingness to swap increases from $0\%$ to 25%, there is a noticeable increase in $\Delta d$. However, from a willingness of $25\%$ onwards, the increase in $\Delta d$ slows down as the willingness of swapping $w\%$ increases.
@@
Moreover, we consider households willingness to swap home in SDGHS under medium tolerance and analyze $\Delta d$ after home swapping under different willingness $w\%$, as shown by the purple line in Figure S8A. The results find that as the $w\%$ increases, $\Delta d$ also increases. When $w=0$, $\Delta d$ is $9.6\%$; when $w>=25$, there is no significant increase in $\Delta d$.
In GHS, we further consider the impact of household members’ selfishness. It is assumed that a household member who experiences an increase in commuting distance will have $s\%$ selfishness to prevent swapping home, even if the average commuting distance of the household decreases. $\Delta d$ after GHS method with different selfishness $s\%$ is shown by the blue line in Figure S8B. The results indicate that $\Delta d$ = $37.5\%$ when households do not swap residence if the commuting distance of a household member increases. As the household members’ selfishness increases, there is a decrease in $\Delta d$. We also consider the impact of household members’ selfishness on SDGHS under medium tolerance and analyze $\Delta d$ after home swapping under different selfishness $s\%$, as shown by the purple line in Figure S8B. The results show that as the household members’ selfishness increases, $\Delta d$ gradually decreases. When s = $100$, $\Delta d$ is $7.9\%$.
VII The Impact of Household Special Needs to Home Swapping
The impacts of educational resources to home swapping. In addition to considering the benefits to commuters, we also considered the impact of educational resources on home swapping. We randomly selected a proportion p of households who refused to relocate, regardless of commuting benefits, and swapped home only for the remaining household (1p1-p). Figure S9A shows the percentage decrease in the average commuting distance after home swapping. We find that, as the proportion p of households who refuses to relocate increases, $\Delta d$ decreases in both GHS and SDGHS (medium tolerance). The results in the figure are averaged over ten independent simulations, It is worth noting that even in the extreme scenario $p$ = $90$ (i.e.,the vast majority of households are immovable), $\Delta d$ still reaches $8.5\%$ in GHS and $2.3\%$ in SDGHS. This suggests that our framework remains effective in reducing congestion even when many households prioritize non-commuting factors. Furthermore, we find that the decrease in $\Delta d$ in SDGHS is significantly slower than that of GHS with increasing p. This phenomenon can be attributed to the pivotal role of ”educational resources” in POI within the context of amenity accessibility in SDGHS. In other words, the results of the simulation can be used to demonstrate again that SDGHS can better portray the impact of the household’s need for a living environment during the process of swapping homes.
Supplementary Figure S9. Impact of residential constraints on commuting distance optimization. (A) Percentage reduction in commuting distance achieved through home-swapping optimization when randomly selecting a percentage of households to remain fixed in their residential locations. (B) Spatial distribution of schools and residential patterns: (top) proportion of families living within increasing region radii (L, 2L…10L) from elementary, middle, high, and all schools, respectively; (bottom) visualization of all schools where circle size scales with student population recorded during the study period. (C) The percentage decrease in average commuting distance under different educational resource constraints, comparing Greedy Home-Swapping (GHS) and Socio-demographic Greedy Home-Swapping (SDGHS) approaches: ’No limitation’ (baseline without educational constraints), ’Part-Fixed’ (fixing residences according to Hebei Province’s 6-18 age population distribution), and ’Distance-Limitation’ (binding households to remain within region radius $L$ of their nearest school). All results represent averages from 10 independent simulations, with SDGHS using medium tolerance parameters.
To better align with real-word constraints, we analyzed the spatial distribution of educational facilities in the target city (Figure S9B). Using the geographic locations of elementary, middle, and high schools, we created concentric regions around each school with radii increasing by $N\times L$ (where L is the city’s average street length) and calculated the proportion of households residing within these regions. The inset map visualizes the city’s hierarchical distribution of educational resources. We then designated households within L meters of any school as ”education-anchored” households, restricting them from participating in residential swaps. As shown in Figure S9C (Distance-Limitation scenario), this constraint affected $19\%$ of households when considering primary and secondary schools. Under this condition, GHS achieved a $40\%$ reduction in commuting distance, while SDGHS still yielded a $10.3\%$ optimization. In addition, according to Hebei Province’s Seventh Population Census, $18\%$ of the population falls within the 6-18 age range (school-attending children). Assuming one student per eligible household, we randomly fixed $18\%$ of urban households to simulate education-driven immobility (Part-Fixed scenario). In this case, GHS reached a $40.3\%$ optimization rate, and SDGHS achieved $10.7\%$. The results presented in the figure have been obtained through the averaging process of 10 independent simulations.
VIII The Reduction of Carbon Emissions
@@
Supplementary Figure S10. The Road carbon emissions obtained based on the empirical data and after RHS and GHS. The top three diagrams show the CO2 emissions on the roads during the morning peak hours between 7-8 AM; the middle three diagrams show the CO2 emissions during the morning peak hours between 8-9 AM; the bottom three diagrams show the CO2 emissions during the evening peak hours between 5-6 PM.
@@
We computed and visualized the carbon emissions of each road segment during the morning peak hours between 7-8 AM, between 8-9 AM, and the evening peak hour between 5-6 PM based on the empirical data and after RHS and GHS, as shown in Figure S10. We found that the carbon emissions under RHS are greater than that based on the empirical data at all time, while those after GHS are significantly less than that of the empirical data. The carbon emissions between 7-8 AM during the morning peak hours are higher than those between 5-6 PM during the evening peak hours. Additionally, after RHS, most road segments emit a large amount of CO2, whereas after GHS, only a few segments have higher emissions, and most road segments emit a small amount of CO2 (road segments emitting 0.4 tonnes of carbon are marked in red, and those emitting between 0.2 and 0.4 tonnes are in yellow, and those emitting less than 0.2 tonnes are in green).
@@
The reduction in carbon emissions after GHS, and SDGHS is shown in Figure S11 for each hour during the peak commuting hours. In the morning peak hours, the average commuting distance after home swapping become shorter for most households. After home swapping, each individual is supposed to arrive at their workplace at the same time before home swapping, which leads to a dispersed departure time (Figure S3). It was found that the reduction in carbon emissions decreases over time. Similarly for the evening peak hours, where each individual is supposed to leave workplace at the same time after home swapping, the shorter commuting distances lead to a gradual increase in carbon emissions reduction over time. Among all the cases, the GHS emissions reduction is the highest, and the emissions reduction of SDGHS is the lowest.
@@
IX Adding New Activity Centers based on the Original City Center
@@
X Impact of Commuting Mode Shift Threshold
@@
While we have assumed 2.5 km to be the shortest distance for commuters to travel by motor vehicles, we also examined other thresholds of assumed minimum distance at $2$, $1.5$, 11, $0.5$, and 0 km, as shown in Figure S15. As the threshold decreases, the proportion of non-motorized vehicles shows a significant decline. Under medium tolerance with SDGHS, 9% of commuters live and work within a very short commuting distance.
@@
In addition, we also explore the impact of the assumed thresholds of the minimum distance of the use of motorized vehicles on the average congestion coefficient and carbon emissions. Figure S16A shows the impact of different values of thresholds on traffic congestion during the 7-8 AM peak hours. As the threshold value decrease, the number of used motor vehicles increases, the occupancy of road sections during peak hours increases, thus increasing congestion. When all commuters travel by motor vehicles, congestion also increases. By considering the three socio-demographic factors, there are fewer non-motorized vehicles, and the change in the minimum travel distance for using motor vehicles does not significantly impact traffic congestion.
Supplementary Figure S16. The impact of the threshold values of the usage of the motorized vehicles. (A) shows the dependence of the average congestion coefficient $I_{t}$ on the threshold values. (B) shows the dependence of the reduction of carbon emissions on the threshold values.